Analysis II Instructor: Russell Brown MWF 9–9:50 Office: POT741 CB341 Phone: 257-3951 Fall 2002 rbrown@uky.edu

Office Hours: MF1-2,W11-12 and by appointment.

Grading: Grades will be based on homework and two longer take home assignments/tests. The homework will be worth 50 points, and each of the longer assignments will be worth 25 points. Collaboration will be encouraged on the homework. The longer assignments should be worked only with the assistance of the instructor. Text: Measure and integral, R. Wheeden and A. Zygmund. The course will cover repeated integration, L^{p} spaces and their basic properties and topics in harmonic analysis, as time permits. Homework: Homework will be assigned and collected regularly. You should endeavor to write out your homework clearly. Use complete sentences. Refer to facts from the text by giving page or result numbers. Note that homework is a substantial fraction of your grade. Be aware that your instructor is old and cranky. Late homework will not be accepted. You may only write on one side of a paper. Leave generous margins. I may use the margins and the back of each sheet for comments. You may also find that your instructor is a bit distracted. Please ask if you to not understand my comments. You may write your homework out by hand. A portion of your homework grade may be based on presentations of homework problems that students give in class. A number of other textbooks cover the material of this course. Three of my favorites are:

• Lebesgue integration on Euclidean spaces, BF Jones.

• Real Analysis, H. Royden

• Real and Complex Analysis, W. Rudin.

Schedule: I have given up trying to predict the future. I hope to cover Chapter 6, Chapter 8, Chapter 9, and additional topics from harmonic analysis which strike our fancy.